# ChemMLLM: Chemical Multimodal Large Language Model

### v1

@@ 3.2 ChemMLLM
Table 5: Results on property2img Task: Hbd, Hba, Rb, and QED (best, 2nd best, *: significantly better).
Chemical molecules are typically encoded in the format of Simplified Molecular Input Line Entry System (SMILES) Weininger (1988), a compact ASCII string, serving as a distinct modality in computational chemistry Anderson et al. (1987). In ChemMLLM, SMILES tokenization is the same as the text. Specifically, SMILES is mapped into a token sequence via Chaemelon Team (2024) text tokenizer trained based on Byte Pair Encoding (BPE) algorithm Sennrich et al. (2015). The process is shown in Figure 1(b).
After training the Mol-VQGAN, ChemMLLM uses it as image tokenizer to align image and text at token level, unifying the training and inference for image and text by maximizing the standard next-token prediction cross-entropy loss:
$\mathcal{L}_{LLM}=$ $\sum_{i=1}^{L}\log p_{\theta}(s_{i}|s_{1},...,s_{i-1})$ (4)
$+\lambda\sum_{k}(\log\sum_{j=1}^{V}\exp(z_{k,j}))^{2},$
where $\lambda$ is a hyper-parameter; in the first term, $L$ is the length of total sequence tokens; $s_{i}\in S=\{S_{I},S_{T}\}$. $S_{I}$ is the image tokens sequence tokenized by the image tokenizer, and $S_{T}$ is the text/SMILES token sequence tokenized by the text tokenizer. The second term is z-loss, a regularization term to mitigate the problem of logit shift in the final softmax and stabilize training Chowdhery et al. (2023), where $z_{k,j}$ denotes the logit in the last layer, $V$ denotes the size of the vocabulary.
Specifically, ChemMLLM adopts Chameleon VQGAN as the image tokenizer/de-tokenizer and Chameleon-7B as the language model. The image tokenizer takes images of 256×256 resolution as input. Simultaneously, text and SMILES information passes through the text tokenizer to be converted to text tokens. The text, SMILES and image tokens are then concatenated to form a unified token sequence to feed into the LLM during training and inference.
3.3 Training
(i) Mol-VQGAN training. The original Chameleon VQGAN is only trained on the natural image dataset and can not discrete and reconstruct molecule images well. So, the first stage focuses on improving VQGAN’s performance in encoding and decoding molecule images. Concretely, we use the well-trained VQGAN (trained on natural images) as the initialization and then fine-tune it on molecule image datasets.
(ii) ChemMLLM Supervised Fine-Tuning Training. In the second stage, we freeze the Mol-VQGAN and only finetune the language model on 5 downstream tasks. We utilize Lumina-mGPT Liu et al. (2024) as training framework to train our ChemMLLM and uses Chameleon-7B as the base model. The weight related to the image tokens in the last layer will first be initialized as zero during finetuning. LLM uses the output of Mol-VQGAN as finetuning data, i.e., the data is first pre-tokenized by Mol-VQGAN and text tokenizer into token sequences and then fed into LLM.
Task Input Output Source # train/test
molecule image captioning (img2caption) image +text text chebi-20 Edwards et al. (2022) mol-instruct Fang et al. (2023) 70K/3K
molecule image property prediction (img2property) image +text text PubChem Kim et al. (2021) 95K/5K
image-to-SMILES conversion (img2smiles) image +text SMILES PubChem Kim et al. (2021) 95K/5K
controllable multi-objective molecule image design (property2img) text image PubChem Kim et al. (2021) 95K/5K
molecule image optimization (img2img) image +text image TDC Huang et al. (2021) 157K/17K
Table 6: Tasks and datasets.
@@ 3.3 Training
Qwen-VL-Chat general text tokenizer, vision encoder $\checkmark$ $\checkmark$ $\times$ $\times$
InternVL-Chat-v1.5 general text tokenizer, vision encoder $\checkmark$ $\checkmark$ $\times$ $\times$
@@ 4 Tasks and Data Curation
(2) Molecule image property prediction (img2property) is an image-to-text task, where models are expected to generate the value of seven different important properties for each molecule image, including molecule weight (MW), Partition Coefficient (P) of a solute between octanol and water (LogP), Topological Polar Surface Area (TPSA), Hydrogen Bond Donor (Hbd), Hydrogen Bond Acceptor (Hba), Rotatable Bond (Rb), and Quantitative Estimate of Drug-likeness (QED). More details for the properties can be found in Appendix G. This image-to-property prediction task evaluates a model’s ability to infer key chemical properties directly from 2D molecular images, enabling researchers to extract actionable insights from molecular images without specialized software, which could accelerate high-throughput screening in drug/material design Lu et al. (2021). Table 12 shows some examples.
@@
(4) Controllable multi-objective molecule image design (property2img) is the inverse problem of molecule image property prediction and is a text-to-image task, where models are expected to generate the image of a molecule conditioned on target properties. It is the core of molecule design Du et al. (2022). The challenge lies in simultaneously optimizing multiple property constraints while maintaining chemical validity. Examples for this task are shown in Table 14.
(5) Molecule image optimization (img2img) is an image-to-image task, where models take a molecular structure with less desirable molecular properties (e.g., LogP) as input and generate a similar molecular structure with more desirable properties while preserving desired chemical properties. It imitates the process of lead optimization, a fundamental problem in drug discovery Huang et al. (2021); Fu et al. (2020). Examples for this task are shown in Table 15.
@@ 4.1 Data Curation
We employ RDKit Landrum et al. (2006) to convert the original SMILES strings into molecular images across all five tasks. We primarily follow the methodology of SketchMol Wang et al. (2025) and ChemVLM Li et al. (2025) for data curation and diversity natural language templates synthesis. The input/output modalities, raw data sources, and sizes of training/test sets for all tasks are shown in Table 6. Further details on data curation are provided in Appendix A.
@@ 5.1 Experimental Setup
Baseline Methods cover both general multimodal LLM and chemical LLM. For general-domain multimodal LLM, we chose Qwen-VL-Chat Bai et al. (2023), InternVL-Chat-v1.5 Chen et al. (2023), LLaVA-v1.5-7B Liu et al. (2023) and GPT-4o OpenAI (2024). For chemical LLM, we chose ChemLLM-7B-Chat Zhang et al. (2024a) and ChemVLM-8B Li et al. (2025). We compare their capabilities in Table 7. Please refer to Appendix D for more descriptions.
Evaluation metrics and implementation details are elaborated in Appendix F and H, respectively. The code is publicly available at https://github.com/bbsbz/ChemMLLM.git.
@@ 5.2 Result
We compare our model with various multimodal LLMs (MLLMs) including Qwen-VL-Chat Bai et al. (2023), InternVL-Chat-v1.5 Chen et al. (2023), LLaVA-v1.5-7B Liu et al. (2023), GPT-4o OpenAI (2024), ChemVLM-8B Li et al. (2025). The evaluation results are shown in Table 1. Our model exhibits strong performance on this task, outperforming all competing MLLM models on all six metrics. An example is shown in Figure 2. Our model generates captions that closely match the ground truth, while Qwen-VL-Chat includes fewer semantically informative details.
@@
GPT-4o can not predict properties from image directly, so in this task we do not compare with GPT-4o. As shown in Table 2 and 3, our model consistently outperforms competing methods across all seven molecular properties, yielding the highest Pearson correlation coefficients alongside the lowest MSE and MAE values. An example is shown in Figure 3. Among the properties predicted, our model has 5 accurate values and 2 close values while Qwen-VL-Chat has 2 close values and 5 inaccurate values.
Model Avg Sim ($\uparrow$) Tani@1.0 ($\uparrow$) valid%($\uparrow$)
Qwen-VL-Chat 0.08 ±$\pm 0.006$ 0.0 ±$\pm 0.0$ 8.2%
InternVL-Chat-v1.5 0.09 ±$\pm 0.003$ 0.0 ±$\pm 0.0$ 20.7%
LLaVA-v1.5-7B 0.05 ±$\pm 0.004$ 0.0 ±$\pm 0.0$ 11.1%
GPT-4o 0.29±$\pm 0.005$ 0.01 ±$\pm 0.004$ 74.5%
ChemVLM-8B 0.55 ±$\pm 0.009$ 0.15 ±$\pm 0.01$ 85.2%
ChemMLLM (ours) 0.75±$\pm 0.009$* 0.49±$\pm 0.01$* 97.1%
Table 8: Results on img2smiles Task. Tanimoto similarities are written as Avg Sim, and Tanimoto@1.0 written as Tani@1.0 (best, 2nd best, *: significantly better).
@@
The evaluation results are shown in Table 8. ChemMLLM performances best in both Tanimoto similarity and Tanimoto@1 metrics. For Tanimoto similarity, ChemMLLM (0.75) surpasses domain-specific model ChemVLM (0.55) by 36.4%. An example is shown in Figure 4. Our model recognizes SMILES from images successfully, while GPT-4o predicts wrong SMILES with low Tanimoto similarity.
@@
Since the GPT-4o API does not perform as well on this task as its web interface, and other MLLMs lack the capability to generate images, we treat this task as a purely text-based problem when evaluating other MLLMs. Specifically, our model is used to generate molecular images, while other MLLMs are tasked with directly generating the corresponding SMILES strings in text form. Given this setting, we also include ChemLLM-Chat-7B Zhang et al. (2024a), a domain-specific chemical language model, in our evaluation. Furthermore, we exclude Qwen-VL-Chat Bai et al. (2023) from comparison on this task, as it fails to generate valid SMILES strings on all test samples. The evaluation results are shown in Table 4 and 5. Our model achieves top-2 or better performance on 90% of the evaluation metrics. Several examples are shown in Figure 5. Our model generates the image molecules with desired properties directly. For more result examples for this task, please refer to Figure 12.
@@
Since the GPT-4o API does not achieve the same level of performance on this task as its web-based interface and other MLLMs lack the ability to generate images, we formulate the evaluation as an image-to-text task for these models. Specifically, our model generates molecular images directly given optimized molecule images, while other MLLMs are required to generate SMILES strings. As GPT-4o cannot directly generate SMILES with optimized LogP from molecular images, we instead treat its evaluation as a text-to-text task, by providing the input SMILES in textual form rather than as images. As shown in Table 9, ChemMLLM achieves the highest increase in LogP, outperforms GPT-4o by 118.9%. Several examples are shown in Figure 6. Our model generates molecule images with higher LogP directly. For more result examples for this task, please refer to Figure 13.
@@
Qwen-VL-Chat -2.0 ±$\pm 0.11$ 0.95±$\pm 0.01$ 1.0±$\pm 0.0$ 4.0%
InternVL-Chat-v1.5 -0.77±$\pm 0.17$ 0.90±$\pm 0.004$ 1.0±$\pm 0.0$ 48.0%
LLaVA-v1.5-7B -0.86±$\pm 0.59$ 0.96±$\pm 0.005$ 1.0±$\pm 0.0$ 37.5%
GPT-4o 1.95±$\pm 0.08$ 0.86±$\pm 0.002$ 1.0±$\pm 0.0$ 99.0%
ChemVLM-8B 0.45 ±$\pm 0.14$ 0.87±$\pm 0.002$ 0.97±$\pm 0.01$ 92.5%
ChemMLLM (ours) 4.2±$\pm 0.44$ 0.88±$\pm 0.001$ 1.0$\pm$0.0 91.0%
Table 9: Results on img2img task (best, 2nd best).
5.3 Ablation Study
We conduct an ablation study on property2img task (see Appendix I) to assess the effects of Mol-VQGAN training and data augmentation (especially image rotation). Results show that both components significantly improve the correlation between generated images and molecular properties, with their combination yielding the best overall performance. This highlights the importance of high-quality visual representations in enhancing multimodal chemical tasks.

### v2

@@ 2 Related Work
Table 2: Results on property2img Task: Hbd, Hba, Rb, and QED (best, 2nd best, *: significantly better (statistically)).
@@
Qwen-VL-Chat (7B) general text tokenizer, vision encoder $\checkmark$ $\checkmark$ $\times$ $\times$
InternVL-Chat-v1.5 (20B) general text tokenizer, vision encoder $\checkmark$ $\checkmark$ $\times$ $\times$
@@
Multimodal LLM (MLLM). With the advancement of large language models and multimodal learning, numerous high-performing multimodal large language models (MLLMs) have recently emerged. Notable MLLMs include GPT-4o (OpenAI 2024), which extends GPT-4V (OpenAI 2023) to understand and generate across different modalities, including text, image, and audio. Emu-3 (Wang et al. 2024) unifies vision understanding and generation via discrete token modeling. Chameleon (Team 2024) aligns modalities at the token level for flexible multimodal generation, while LlamaGen (Sun et al. 2024) treats images as language-like sequences for scalable autoregressive image generation. Lumina-mGPT (Liu et al. 2024) trains a family of models to generate flexible photorealistic images from text descriptions based on Chameleon. Transfusion (Zhou et al. 2024) and Show-o (Xie et al. 2024) combine diffusion and transformer for multimodal understanding and generation. For vision understanding, well-known models include LLaVA (Liu et al. 2023), BLIP-2 (Li et al. 2023), Qwen-VL-Chat (Bai et al. 2023), InternVL-Chat (Chen et al. 2023), MiniGPT-4 (Zhu et al. 2023), Gemini (Team et al. 2023), Flamingo (Alayrac et al. 2022) and OpenFlamingo (Awadalla et al. 2023). Although current MLLMs perform well across modalities, these models struggle with chemical tasks due to a misalignment between general and domain-specific knowledge.
Chemical LLM. MLLMs have also exhibited strong potential in addressing chemistry-related tasks, particularly in bridging the modality gap between textual descriptions and molecular representations. Concretely, Instruct-Mol (Cao et al. 2023) and MV-Mol (Luo et al. 2024) utilize LLaVA’s architecture (Liu et al. 2023) and Q-former (Li et al. 2023) to align molecular structure and text modality, respectively. ChemLLM (Zhang et al. 2024a) uses a high-quality chemical dataset to fine-tune InternLM2 (Cai et al. 2024). ChemVLM (Li et al. 2025) extends ChemLLM (Zhang et al. 2024a) to understand images by adopting a projector-based method to align vision information and text information. UniMoT (Zhang et al. 2024b) uses a molecule tokenizer to align the graph modality molecule with text. Despite progress in chemical tasks, existing models can not generate molecular images, limiting their utility in more intuitive, visual forms of interaction. To fill this gap, we propose ChemMLLM, a unified framework that understands and generates molecules in text, SMILES, and image formats.
3 Method
Overview. Our framework uses an image tokenizer to transfer images into discrete tokens, aligning with texts and molecule’s SMILES strings at the token level, known as “Image Tokenizer-LLM-Image De-tokenizer” architecture (Team 2024). We first discuss the molecule image tokenizer, and then describe how to combine images, texts and SMILES strings in Multimodal LLM, followed by the training strategy. The whole pipeline is shown in Figure 2. Also, we list key mathematical notations in Table 8 in Appendix B.
3.1 Mol-VQGAN for Molecule Image Generation
[…]
$\hat{x}=G(z_{q})=G(\textbf{q}(E(x))).$ (2)
The discretization and reconstruction process is illustrated in Figure 2(a). The Mol-VQGAN’s objective function is:
$\min_{E,G,Z}\max_{D}\big{[}\mathcal{L}_{vqvae}(E,G,Z)$ (3)
$+\lambda_{1}\mathcal{L}_{perceptual}(E,G,Z)+\lambda_{2}\mathcal{L}_{GAN}(\{E,G,Z\},D)\big{]},$
where (1) the first term $\mathcal{L}_{vqvae}(E,G,Z)=||x-G(\hat{z}-sg(\hat{z}-z_{q}))||_{2}^{2}+||sg[\hat{z}]-z_{q}||^{2}_{2}+||sg[z_{q}]-\hat{z}||^{2}_{2}$ is VQVAE loss, where $\mathcal{L}_{rec}=||x-G(\hat{z}-sg(\hat{z}-z_{q}))||_{2}^{2}$ is reconstruction loss. Since the quantization process is non-differentiable, a stop-gradient operation $sg[\cdot]$ is used so that the forward pass operates on quantized vectors, whereas the backward pass leverages continuous vectors for gradient computation (Van Den Oord, Vinyals et al. 2017); (2) the second term $\mathcal{L}_{perceptual}(E,G,Z)=||P(x)-P(G(\hat{z}-sg(\hat{z}-z_{q})))||_{2}^{2}$ is perceptual loss, $P$ denotes a perceptual model like Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al. 2018), which is used to extract the high-level semantic features; (3) the third term $\mathcal{L}_{GAN}(\{E,G,Z\},D)=\log D(x)+\log(1-D(\hat{z}-sg(\hat{z}-z_{q})))$ is GAN loss, $D$ is a patch-based discriminator (Isola et al. 2017) aiming to differentiate original and reconstructed images. $\lambda_{1}$ is a hyperparameter and $\lambda_{2}$ is an adaptive weight computed dynamically to balance the weight of $\mathcal{L}_{GAN}$ and stabilize training, following VQGAN (Esser, Rombach, and Ommer 2021).
3.2 ChemMLLM
Chemical molecules are typically encoded in the format of Simplified Molecular Input Line Entry System (SMILES) (Weininger 1988), a compact ASCII string, serving as a distinct modality in computational chemistry (Anderson, Veith, and Weininger 1987). In ChemMLLM, SMILES tokenization is the same as the text. Specifically, SMILES is mapped into a token sequence via Chameleon (Team 2024) text tokenizer trained based on Byte Pair Encoding (BPE) algorithm (Sennrich, Haddow, and Birch 2015). The process is shown in Figure 2(b).
ChemMLLM uses a well-trained Mol-VQGANas image tokenizer to align image and text at token level, unifying the training and inference for image and text by maximizing the standard next-token prediction cross-entropy loss:
$\mathcal{L}=\sum_{i=1}^{L}\log p_{\theta}(s_{i}|s_{1},.,s_{i-1})+\lambda\sum_{k}(\log\sum_{j=1}^{V}\exp(z_{k,j}))^{2},$ (4)
where $\lambda$ is a hyper-parameter; in the first term, $L$ is the length of total sequence tokens; $s_{i}\in S=\{S_{I},S_{T}\}$. $S_{I}$ is the image tokens sequence tokenized by the image tokenizer, and $S_{T}$ is the text/SMILES token sequence tokenized by the text tokenizer. The second term is z-loss, a regularization term to mitigate the problem of logit shift in the final softmax and stabilize training (Chowdhery et al. 2023), where $z_{k,j}$ denotes the logit in the last layer, $V$ denotes the size of the vocabulary.
Specifically, ChemMLLM adopts Chameleon VQGAN as the image tokenizer/de-tokenizer and Chameleon-7B as the language model. The image tokenizer takes images of 256×256 resolution as input. Simultaneously, text and SMILES information pass through the text tokenizer to be converted to text tokens. The text, SMILES and image tokens are then concatenated to form a unified token sequence to feed into the LLM during training and inference.
3.3 Training
(i) Mol-VQGAN training. The original Chameleon VQGAN is only trained on the natural image dataset and can not discretize and reconstruct molecule images well. So, the first stage focuses on improving VQGAN’s performance in encoding and decoding molecule images. Concretely, we use the well-trained VQGAN (trained on natural images) as the initialization and then fine-tune it on molecule image datasets.
(ii) ChemMLLM Supervised Fine-Tuning (SFT). In the second stage, we freeze the Mol-VQGAN and only finetune the language model on 5 downstream tasks. We utilize Lumina-mGPT (Liu et al. 2024) as training framework to train our ChemMLLM and use Chameleon-7B as the base model. The weights related to the image tokens in the last layer will first be initialized as zero during finetuning. LLM uses the output of Mol-VQGAN as finetuning data, i.e., the data is first pre-tokenized by Mol-VQGAN and text tokenizer into token sequences and then fed into LLM.
Task Input Output Source # train/test
molecule image
captioning (img2caption)
image
+text
text
chebi-20 (Edwards et al. 2022)
Mol-Instructions (Fang et al. 2023)
70K/3K
molecule image property
prediction (img2property)
image
+text
text PubChem (Kim et al. 2021) 200K/5K
image-to-SMILES
conversion (img2smiles)
image
+text
SMILES PubChem (Kim et al. 2021) 200K/5K
controllable multi-objective
molecule image design
text image PubChem (Kim et al. 2021) 200K/5K
molecule image
optimization (img2img)
image
+text
image TDC (Jin et al. 2018) 157K/17K
Table 4: Tasks and datasets.
@@ 4 Tasks and Data Curation
(2) Molecule image property prediction (img2property) is an image-to-text task, where models are expected to generate the value of seven different important properties for each molecule image, including molecule weight (MW), Partition Coefficient (P) of a solute between octanol and water (LogP), Topological Polar Surface Area (TPSA), Hydrogen Bond Donor (Hbd), Hydrogen Bond Acceptor (Hba), Rotatable Bond (Rb), and Quantitative Estimate of Drug-likeness (QED). More details for the properties can be found in Appendix G. This image-to-property prediction task evaluates a model’s ability to infer key chemical properties directly from 2D molecular images, enabling researchers to extract actionable insights from molecular images without specialized software, which could accelerate high-throughput screening in drug/material design (Lu et al. 2021). Table 10 shows some examples.
@@
(4) Controllable multi-objective molecule image design (property2img) is the inverse problem of molecule image property prediction and is a text-to-image task, where models are expected to generate the image of a molecule conditioned on target properties. It is the core of molecule design (Du et al. 2022). The challenge lies in simultaneously optimizing multiple property constraints while maintaining chemical validity. Examples for this task are shown in Table 12.
(5) Molecule image optimization (img2img) is an image-to-image task, where models take a molecular structure with less desirable molecular properties (e.g., LogP) as input and generate a similar molecular structure with more desirable properties while preserving desired chemical properties. It imitates the process of lead optimization, a fundamental problem in drug discovery (Huang et al. 2021; Fu, Xiao, and Sun 2020). Examples for this task are shown in Table 13.
@@ 4.1 Data Curation
We employ RDKit (Landrum et al. 2006) to convert the original SMILES strings into molecular images across all five tasks. We primarily follow the methodology of SketchMol (Wang et al. 2025) and ChemVLM (Li et al. 2025) for data curation and diversity natural language templates synthesis. The input/output modalities, raw data sources, and sizes of training/test sets for all tasks are shown in Table 4. Further details on data curation are provided in Appendix A.
@@ 5.1 Experimental Setup
Baseline methods cover (i) general multimodal LLM such as Qwen-VL-Chat (Bai et al. 2023), InternVL-Chat-v1.5 (Chen et al. 2023), LLaVA-v1.5-7B (Liu et al. 2023) and GPT-4o (OpenAI 2024); (ii) chemical LLM such as ChemLLM-7B-Chat (Zhang et al. 2024a) and ChemVLM-8B (Li et al. 2025); (iii) task-specific models such as sequence-to-sequence model (Cho et al. 2014), Chemception (Goh et al. 2017) (convolutional neural network), Chemprop (Heid et al. 2023) (graph neural network), Chemformer (Irwin et al. 2022), MolScribe (Qian et al. 2023) and Decimer (Rajan, Zielesny, and Steinbeck 2020). We compare their capabilities in Table 3. Please refer to Appendix D for more descriptions.
Evaluation metrics and implementation details are elaborated in Appendix F and  H. The code is publicly available at https://github.com/bbsbz/ChemMLLM.git.
@@ 5.2 Result
Overview. As shown in Figure 3, our ChemMLLM outperforms the strongest baseline in all five tasks consistently.
Figure 3: Performance of ChemMLLM-7B, ChemMLLM-34B and the best baseline on five tasks. Mean Pearson means the mean value of Pearson correlation of seven properties; Avg Sim means Tanimoto similarities; Normalized $\Delta$LogP means normalized (i.e., divided by the maximum value) Increased LogP.
@@
We compare our model with various multimodal LLMs (MLLMs), including Qwen-VL-Chat (Bai et al. 2023), InternVL-Chat-v1.5 (Chen et al. 2023), LLaVA-v1.5-7B (Liu et al. 2023), GPT-4o (OpenAI 2024), ChemVLM-8B (Li et al. 2025). Due to space limitations, experimental results are provided in Appendix L.1. As shown in Table 16 (Appendix), our model exhibits strong performance on this task, outperforming all competing MLLM models on all six metrics. An example is shown in Figure 4. Our model generates captions that closely match the ground truth, while Qwen-VL-Chat includes fewer semantically informative details.
@@
For this task, we do not compare with GPT-4o because it cannot predict properties from images directly, and we compare with task-specific models, including Chemception (Goh et al. 2017) and Chemprop (Heid et al. 2023). Due to space limitation, the experimental results are presented in Appendix 16. As shown in Table 14 and 15, our model consistently outperforms MLLMs across all seven molecular properties, yielding the highest Pearson correlation coefficients alongside the lowest MSE and MAE values. While ChemMLLM’s performance is behind specialized models like Chemprop, it offers greater versatility in handling a wider range of tasks. An example is shown in Figure 5. Among the properties predicted, our model has 5 accurate values and 2 close values, while Qwen-VL-Chat has 2 close values and 5 inaccurate values.
Model Avg Sim ($\uparrow$) Acc ($\uparrow$) valid%($\uparrow$)
MolScribe 0.98 ±0.002$\pm$ 0.002± 0.002 0.66 ±0.01$\pm$ 0.01± 0.01 96.9%
Decimer 0.97 ±0.002$\pm$ 0.002± 0.002 0.78 ±0.01$\pm$ 0.01± 0.01 99.9%
Qwen-VL-Chat (7B) 0.08 ±0.006$\pm$ 0.006± 0.006 0.0 ±0.0$\pm$ 0.0± 0.0 8.2%
InternVL-Chat-v1.5 (20B) 0.09 ±0.003$\pm$ 0.003± 0.003 0.0 ±0.0$\pm$ 0.0± 0.0 20.7%
LLaVA-v1.5-7B 0.05 ±0.004$\pm$ 0.004± 0.004 0.0 ±0.0$\pm$ 0.0± 0.0 11.1%
GPT-4o 0.29±0.005$\pm$ 0.005± 0.005 0.01 ±0.004$\pm$ 0.004± 0.004 74.5%
ChemVLM-8B 0.55 ±0.009$\pm$ 0.009± 0.009 0.11 ±0.01$\pm$ 0.01± 0.01 85.2%
ChemMLLM-7B (ours) 0.75±0.009$\pm$ 0.009± 0.009* 0.39±0.01$\pm$ 0.01± 0.01* 97.1%
ChemMLLM-34B (ours) 0.87±0.007$\pm$ 0.007± 0.007* 0.56±0.01$\pm$ 0.01± 0.01* 97.2%
Table 5: Results on img2smiles Task. “Avg Sim” means Tanimoto similarities, and “Acc” means Accuracy (best, 2nd best, *: significantly better). Task-specific models Decimer (Rajan, Zielesny, and Steinbeck 2020) and MolScribe (Qian et al. 2023) leverage CNN+Transformer and CNN+GNN architectures, respectively.
@@
In this task, we compare our model with MLLMs and task-specific models including MolScribe (Qian et al. 2023) and Decimer (Rajan, Zielesny, and Steinbeck 2020). The evaluation results are shown in Table 5. ChemMLLM achieves the best performance in both Tanimoto similarity and Accuracy metrics among all MLLMs. Despite exhibiting lower performance than MolScribe and Decimer, our model is capable of handling a broader range of downstream tasks beyond img2smiles. For Tanimoto similarity, ChemMLLM (0.87) surpasses domain-specific model ChemVLM (0.55) by 58.18%. An example is shown in Figure 6. Our model recognizes SMILES from images successfully, while GPT-4o predicts wrong SMILES with low Tanimoto similarity.
@@
Since the GPT-4o API does not perform as well on this task as its web interface, and other MLLMs lack the capability to generate images, we treat this task as a purely text-based problem when evaluating other MLLMs. Specifically, our model is asked to generate molecular images, while other MLLMs are tasked with directly generating the corresponding SMILES strings in text form. Given this setting, we also include ChemLLM-Chat-7B (Zhang et al. 2024a), a domain-specific chemical language model, in our evaluation. Furthermore, we exclude Qwen-VL-Chat (Bai et al. 2023) from comparison on this task, as it fails to generate valid SMILES strings on all test samples. The evaluation results are shown in Table 1 and 2. Our model achieves top-2 or better performance on all evaluation metrics. Several examples are shown in Figure 7. Our model generates the molecule images with desired properties directly. For more result examples for this task, please refer to Figure 14.
@@
Since GPT-4o’s API underperforms its web interface on this task and other MLLMs lack image-generation capability, we evaluate them as image-to-text models (generating SMILES from images). For GPT-4o, which cannot produce optimized SMILES directly from images, we adopt a text-to-text approach, providing input SMILES as text instead. We also compare ChemMLLM with task-specific models including sequence-to-sequence (Cho et al. 2014) and Chemformer (Irwin et al. 2022). As shown in Table 6, ChemMLLM achieves the highest increase in LogP, outperforms GPT-4o by 116.75%. Several examples are shown in Figure 8. Our model generates molecule images with higher LogP directly. For more result examples for this task, please refer to Figure 15.
@@
Seq2Seq 1.95 ±0.12$\pm$ 0.12± 0.12 0.79±0.007$\pm$ 0.007± 0.007 1.0±0.0$\pm$ 0.0± 0.0 80.5%
Chemformer 3.03 ±0.08$\pm$ 0.08± 0.08 0.85 ±0.002$\pm$ 0.002± 0.002 0.93 ±0.01$\pm$ 0.01± 0.01 100%
Qwen-VL-Chat (7B) 1.50 ±0.93$\pm$ 0.93± 0.93 0.95±0.01$\pm$ 0.01± 0.01 1.0±0.0$\pm$ 0.0± 0.0 4.0%
InternVL-Chat-v1.5 (20B) 0.77±0.23$\pm$ 0.23± 0.23 0.90±0.004$\pm$ 0.004± 0.004 1.0±0.0$\pm$ 0.0± 0.0 48.0%
LLaVA-v1.5-7B 1.72±1.53$\pm$ 1.53± 1.53 0.96±0.005$\pm$ 0.005± 0.005 1.0±0.0$\pm$ 0.0± 0.0 37.5%
GPT-4o 1.97±0.07$\pm$ 0.07± 0.07 0.86±0.002$\pm$ 0.002± 0.002 1.0±0.0$\pm$ 0.0± 0.0 99.0%
ChemVLM-8B 0.67 ±0.13$\pm$ 0.13± 0.13 0.87±0.002$\pm$ 0.002± 0.002 0.97±0.01$\pm$ 0.01± 0.01 92.5%
ChemMLLM-7B (ours) 4.27±0.44$\pm$ 0.44± 0.44* 0.88±0.001$\pm$ 0.001± 0.001 1.0$\pm$0.0 91.0%
ChemMLLM-34B (ours) 4.07±0.38$\pm$ 0.38± 0.38* 0.88±0.001$\pm$ 0.001± 0.001 1.0$\pm$0.0 99.0%
Table 6: Results on img2img task (best, 2nd best, *: significantly better). Task-specific models Seq2seq (Cho et al. 2014) and Chemformer (Irwin et al. 2022) leverage Gated Recurrent Unit (GRU) and Bidirectional and Auto-Regressive Transformers (BART) (Lewis et al. 2019) architectures, respectively.
Ablation Study. We conduct an ablation study on property2img task (Appendix I) to assess the effects of two-stage training (Mol-VQGAN training and ChemMLLM SFT) and data augmentation (especially image rotation). Results show that both components significantly improve the correlation between generated images and molecular properties, with their combination yielding the best overall performance. This highlights the importance of high-quality visual representations in enhancing multimodal chemical tasks.
We also empirically investigate the influence of the hyperparameter $\lambda$ in Eq. (4) (see Appendix J.1), and find that the impact of $\lambda$ in z-loss is marginal.